Find the product of 32 and 46. Now reverse the digits and find the product of 23 and 64. The products are the same. Does this happen with any pair of two-digit numbers? Find two other pairs of two-digit numbers (with different digits) that have this property.
Explain your thinking in solving this problem and also include any mathematical work.

Answer by checkley77(12844)

32*46=1,472
23*64=1,472
Technique:
Pick a number:
41 - multiply by 2=82 - now reverse the digits=28
41*28=1,148
14*82=1,148